OpenStudy (anonymous):

please help? what is the equation in standard form, of a parabola that contains the following points? (-2,18), (0,4), (4,24) y=-2x^2+3x-4 y=2x^2-3x+4 y=-3x^2+2x+4 y=_4x^2-3x-2

OpenStudy (anonymous):

the last one should be -4, not _4

OpenStudy (anonymous):

anybody?

OpenStudy (tkhunny):

Have you considered substitution and the solution of the resulting simultaneous equations?

OpenStudy (anonymous):

how would I do that?

OpenStudy (tkhunny):

From the given choices, it's pretty clear that the equation will look like this: y = ax^2 + bx + c Right?

OpenStudy (anonymous):

oh wait I just found the example in my book. lets see here...

OpenStudy (tkhunny):

Don't miss the fact that the GAVE US the y-intercept. This leads immediately to: a(0)^2 + b(0) + c = 4 c = 4 Now, we have this: y = ax^2 + bx + 4 And two more points to find a and b.

OpenStudy (anonymous):

I believe it would be y=2x^2-3x+4, right?

OpenStudy (tkhunny):

Did you just substitute the points into the given equations? BTW Once we have c = 4, we can discard the first and the last.

OpenStudy (anonymous):

im solving it how it showed me in the book on paper, and the answer I posted is what I got after substituting the points into the equation.

OpenStudy (tkhunny):

Unique answers don't care how you find them. Good work.

OpenStudy (anonymous):

sweet. thanks for your help